OpenStudy (anonymous):

y + 2x = 5;x = -1,0,3 find the value of y for each value of x

OpenStudy (anonymous):

y=7,5,-1 respectively

OpenStudy (osanseviero):

You need to do some basic arithmetic. The problem gives you a value for x. You just need to put the value of x. First, you need to know this: y + 2x = 5 So: y + 2x - 2x = 5 -2x y = 5 - 2x If x = -1: y = 5 - 2 (-1) = 5 + 2 = 7 Try the next one yourself

OpenStudy (anonymous):

so the next one is 3x - 5y =9; x =-1,0,1

OpenStudy (anonymous):

so whats the first step

OpenStudy (osanseviero):

First of all, you want to now the value of y. So the first step is to have y alone